OpenStudy (anonymous):

how do I simplify 2cos2ysin2y

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

You apply the double angle formula for Sine which states: \[\sin(2x)=2\sin(x)\cos(x)\] From here You can work backwards and say: Let x=2y. Or: \[2\sin(2y)\cos(2y)=2\sin(x)\cos(x)=\sin(2x)\] But since you let x=2y it becomes: \[\sin(4y)\]

OpenStudy (anonymous):

That helps - thank you very much!

OpenStudy (anonymous):

No problem :)